# Planning Agents on an Ego-Trip: Leveraging Hybrid Ego-Graph Ensembles for Improved Tool Retrieval in Enterprise Task Planning

### v1

@@ 6 Limitations
While our approach improves tool retrieval performance for several complex query types, it has notable limitations. First, the effectiveness depends heavily on the quality and completeness of the underlying knowledge graph. Errors in triple extraction or missing tool relationships can weaken results considerably. Second, the framework may be less robust in domains with highly heterogeneous or sparsely described tools, reducing node matching accuracy.
@@
Finally, while we demonstrate strong results in our enterprise task domain, broader evaluation is needed to assess generalizability across different domains and tool ecosystems.
@@ 7 Conclusion and Future Work
The task of efficiently exploring available tools—crucial for effective task decomposition during agentic planning—remains challenging for LLM-powered systems, particularly in enterprise environments with numerous tools that have complex and often undocumented interdependencies. Our research was guided by the hypothesis that explicitly modeling these relationships in a graph structure would enhance tool retrieval effectiveness. To address the scarcity of complex queries in existing datasets, we propose a synthetic query generation pipeline that models tool dependencies through parameter-level connections, enabling the generation of realistic, multi-step queries. We presented a systematic approach to transform enterprise tools into a coherent graph representation and introduced a novel Ensemble of Ego-Graphs (EEG) retrieval framework that outperforms traditional baselines. Our results as shown in Table 3 as well as appendix B.2 support our hypothesis and establish a promising direction for improving tool retrieval in complex enterprise environments. Future research directions addressing some of the shortcomings we have identified include:
@@ 8 GenAI Usage Disclosure
We employed ChatGPT and Claude to assist in rephrasing certain sections of the paper for improved clarity. All core content, including research design, data analysis, and result interpretation, was conducted with- out the aid of generative AI tools.
@@ Appendix A Complete Example
Figure 4: Relevant subgraph to help illustrate the example
Appendix B More Experimental Results
B.1 Representative Synthetic Query Examples
Table 4 presents detailed representative examples for queries generated for each query type.
Query Type Example
B.2 CompleteRecall@10 per Query Category

### v2

@@ 6 Limitations
While our approach improves tool retrieval performance for several complex query types, it has notable limitations. First, its effectiveness depends heavily on the quality and completeness of the underlying knowledge graph. Errors in triple extraction or missing tool relationships can considerably weaken results and create a lower bound on achievable performance, since inaccuracies in the extraction phase may propagate through later stages. Second, the framework may be less robust in domains with highly heterogeneous or sparsely described tools, reducing node-matching accuracy.
@@
Finally, while we demonstrate strong results in our enterprise task domain, broader evaluation is needed to assess generalization across different domains and tool ecosystems.
@@ 7 Conclusion and Future Work
The task of efficiently exploring available tools, which is crucial for effective task decomposition during agentic planning, remains challenging for LLM-powered systems, particularly in enterprise environments with numerous tools that have complex and often undocumented interdependencies. Our research was guided by the hypothesis that explicitly modeling these relationships in a graph structure would enhance tool retrieval effectiveness. To address the scarcity of complex queries in existing datasets, we propose a synthetic query generation pipeline that models tool dependencies through parameter-level connections, enabling the generation of realistic, multi-step queries.
We present a systematic approach to transform enterprise tools into a coherent graph representation and introduc a novel Ensemble of Ego-Graphs (EEG) retrieval framework that outperforms traditional baselines. Our results as shown in Table 4 support our hypothesis and establish a promising direction for improving tool retrieval in complex enterprise environments.
Future research directions addressing some of the shortcomings we have identified include:
@@
  • Adding additional dimensions to our dataset to go beyond the current defined classes and better capture the real-world variability and messiness of user queries.
@@
  • Evaluating and optimizing graph-retrieval latency to make it comparable with current retrieval mechanisms
  • Introducing query augmentation strategies to inject realistic linguistic variability, such as ambiguity, underspecification, and incomplete phrasing to improve generalizability to real-world enterprise queries
  • Incorporating inter-annotator agreement measures (e.g., Cohen’s Kappa) to evaluate consistency among experts and between experts and the automated validation pipeline to strengthen the reliability of evaluation outcomes
@@ 8 GenAI Usage Disclosure
We employed ChatGPT and Claude to assist in rephrasing certain sections of the paper for improved clarity. All core content, including research design, data analysis, and result interpretation, was conducted without the aid of generative AI tools.
@@ Appendix A Complete Example
Figure 5: Subgraph to help illustrate the example in Appendix A
Appendix B CompleteRecall@10 per Query Category